# Generation of structure-guided pMHC-I libraries using Diffusion Models

### v1

@@ 4 Discussion and limitations
Our study addresses critical limitations in pMHC-I binding prediction by leveraging diffusion models conditioned on atomic-level interactions, effectively avoiding biases inherent in mass-spectrometry and binding assay data. The structure-aware generative method introduced here challenges current models, highlighting their inability to recognize structurally valid, novel peptides that are out-of-distribution for their experimental training datasets.
@@
Our workflow and generated library of $10^{5}$ peptides serves both as a benchmark for future predictive models and as a method to expose data biases. While validation of our designed peptides primarily relied on the unbiased EpiScan dataset, which is limited to four alleles, our method is generalizable to a broader range of HLA types. Additionally, incorporating TCR-binding predictions will extend our methodology’s relevance to comprehensive immunotherapy design. Future research should prioritize broader experimental work and integration of TCR interactions to fully capture the complexity of immune responses.
Appendix A Supplemental Figures
Figure S1: JS divergence matrix between 9-mer PWMs extracted from high-confidence (pLDDT $\geq$ 0.8) peptides bound in each crystal structure. Axes list PDB IDs; side-bars encode the restricting HLA class I allele.
Figure S2: JS Divergence of the Structurally-aware peptides against the Publicly available peptides
@@ Appendix A Supplemental Figures
Figure S3: Comparison of EpiScan and this study’s results for peptide-MHC interactions for HLA-A*02:01.
@@
Figure S4: Comparison of EpiScan and this study’s results for peptide-MHC interactions for HLA-B*57:01.
Figure S5: Scatterplots of the raw scores for each of the predictors across the confidence levels of each peptide.
Figure S6: Sequence logos of structure-guided peptide libraries across HLA alleles and control conditions. Each row represents a different HLA allele, with three columns showing: (1) Anchoring - peptides generated using our structure-guided diffusion method with hot-spot residues anchored to the MHC binding pocket; (2) No anchoring (pLDDT $>$ 0.8) - control peptides where hot-spot residues are anchored to random MHC surface positions outside the binding pocket, filtered for high structural confidence; (3) No anchoring (no pLDDT filter) - control peptides with random surface anchoring without structural confidence filtering. The random surface anchoring serves as a negative control to demonstrate that biologically relevant binding motifs emerge only when hot-spot residues are constrained to the actual binding pocket.
(continued) Additional alleles and control conditions (see Fig. S6 for full legend).
Figure S7: Sequence novelty versus predicted binding affinity for different peptide generation methods. Euclidean distance to public datasets measures sequence novelty relative to known epitope position weight matrices, while ESMCBA log10 IC50 predictions indicate predicted binding strength. LM (Language Model) refers to peptides generated by autoregressive sampling from ESM2, showing high novelty but poor predicted binding. Anchor represents anchor-preserved permutation controls with low novelty but variable binding predictions. RFdiffusion peptides uniquely combine high sequence novelty with favorable predicted binding affinities, demonstrating the method’s ability to explore functionally relevant regions of sequence space beyond current database coverage.
Figure S8: UMAP visualization of ESMCBA embeddings for HLA-A*02:01 peptides across different generation methods and scoring metrics. The first row shows the key peptide categories for sequence space analysis: Experimental IC50 Measurement (experimentally validated epitopes), Anchor-Preserved Mutagenesis Baseline (control peptides with conserved anchor residues but randomized non-anchor positions), and Global Sequence Sampling Control (random peptides generated by autoregressive ESM2 sampling). These correspond to the three main categories discussed in the sequence novelty analysis. The remaining panels show various computational predictions and structural confidence scores colored on the same embedding space.

### v2

@@ 4 Discussion and limitations
Our study addresses critical limitations in pMHC-I binding prediction by leveraging diffusion models conditioned on atomic-level interactions, effectively avoiding biases inherent in mass-spectrometry and binding assay data. The structure-aware generative method introduced here challenges current models, highlighting their inability to recognize structurally valid, novel peptides that are out-of-distribution for their experimental training datasets. In particular, our approach recapitulates canonical anchoring preferences observed in unbiased datasets such as EpiScan (for example, recovery of tryptophan, valine, and leucine anchors) and shows that these residues systematically align with higher predicted structural confidence (iPTM and pLDDT). This anchor recovery provides confidence that our approach captures allele-specific motifs despite not being trained directly on binding data. At the same time, our methodology expands the explored sequence space, generating diverse peptides that probe regions of the epitope landscape underrepresented in current datasets.
Several limitations should be acknowledged. First, we observed a non-trivial enrichment of proline across alleles. This likely reflects the well-documented tendency of diffusion-based generative models to rigidify local structures by favoring proline, a behavior also noted by the authors of RFdiffusion in other contexts. While proline enrichment may provide stabilizing biophysical effects in some cases, its consistent appearance across unrelated alleles suggests that it can also represent a technical artifact of the generative process rather than a true biological signal. Second, our positional variance analysis revealed additional biases in amino-acid usage that were relatively uniform across alleles. Such across-allele uniformity may indicate model-driven design constraints that override allele-specific preferences, highlighting the importance of correcting for background biases when interpreting sequence logos or enrichment statistics. Third, although our approach successfully recapitulates canonical anchors and these residues were associated with systematically higher structural confidence, structural proxies remain imperfect surrogates for physical binding. The iPTM score, which measures AlphaFold2’s confidence in interface packing, and the per-residue pLDDT metric are both useful correlates, but they cannot directly prove biochemical stability or immunogenicity. Finally, by design our approach expands into sequence space that lies outside the distribution of known epitopes. This novelty is a strength, as it enables exploration of under-sampled regions of the binding landscape, but it also means that our generated peptides must ultimately be validated experimentally. Without functional testing, their true binding capacity and immunological relevance remain unknown.
@@
Overall, our workflow and generated library provide both a benchmark for predictive models and a tool to uncover data- and model-driven biases. While our validation relied primarily on the unbiased EpiScan dataset, which is limited to four alleles, our framework is generalizable to a broader range of HLA types. Looking forward, incorporating TCR-binding predictions and extending experimental validation across additional alleles will broaden the relevance of this methodology for comprehensive immunotherapy design. In doing so, our approach establishes a path toward expanding the epitope landscape beyond current data limitations while maintaining biological plausibility through structural constraints.
5 Data and Code Availability
Data and code for all the analyses can be found in the Github repository: https://github.com/sermare/struct-mhc-dev.
6 Acknowledgements
We thank Dr. Peter Bruno for valuable comments and help with the EpiScan analysis.
Appendix A Supplemental Tables
HLA Allele # Structures PDB IDs
A0201 15 7KGP, 7KGQ, 7LG2, 7MJ9, 7MKB, 7MJ6, 3TO2, 7KGO, 7KGS, 7MJ8, 7LG3, 7MJ7, 7KGT, 7SA2, 3OXS
A0206 1 3OXR
A0207 1 3OXS
A2 6 6UJQ, 6O4Z, 6O4Y, 6UJO, 6O51, 6O53
A2402 7 7JYW, 7EJN, 8SBK, 7EJL, 7JYV, 7EJM, 8HN4
A6 1 4HX1
B1301 1 7YG3
B1501 5 6UZS, 7XF3, 6VB3, 6UZQ, 6UZP
B1502 11 6VB6, 6VIU, 6VB7, 6UZM, 6UZO, 6VB1, 6UZN, 6VB0, 6VB4, 6VB2, 6VB5
B1801 4 8ROP, 8RNH, 8ROO, 8RNG
B2705 7 6Y28, 6VQE, 6VQD, 6VQ2, 6VPZ, 6Y2A, 6Y26
B2706 1 5DEG
B3501 7 8EMK, 8EMF, 7SIG, 8EMG, 8EMI, 8EMJ, 4PRA
B3503 1 7SIH
B3505 1 7SIF
B3508 13 4PRD, 3VFP, 3VFR, 3VFN, 3VFO, 3VFU, 7KGO, 3VFS, 3VFM, 3VFT, 4PRE, 3VFW, 3VFV, 4PRB
B4001 1 6IEX
B4403 2 3KPN, 3KPO
B4405 2 6MTL, 1SYV
B5301 2 7R7W, 7R7V
B5701 3 6BXQ, 7R7Y, 6V2O
B5703 2 5VVP, 5VWD
B5801 2 7X00, 7WZZ
B5803 1 5VWF
B8 3 8E8I, 8EC5, 8E13
C1202 1 9L48
C1203 1 9L49
N0101 1 6K7T
Table 1: HLA alleles with associated PDB structures
Appendix B Supplemental Figures
Figure S1: Jensen-Shannon distance matrix between 9-mer PWMs extracted from high-confidence (pLDDT $\geq$ 0.7) from the structure-guided generated peptides.
Figure S2: Amino acid frequencies across positions in 9-mer structure-guided peptides from 27 alleles.
@@ Appendix B Supplemental Figures
Figure S5: Sequence logos of 9-mer peptides grouped by allele, normalized against positional background frequencies. The height of each letter reflects relative information content.
Figure S6: UMAP visualization of ESMCBA embeddings for HLA-A*02:01 peptides across classes and scoring metrics. The first row shows the key peptide categories for sequence space analysis: Experimental IC50 Measurement (experimentally validated epitopes), Anchor-Preserved Mutagenesis Baseline (control peptides with conserved anchor residues but randomized non-anchor positions), and Global Sequence Sampling Control (random peptides generated by autoregressive ESM2 sampling).